price =10 birr and tc=120+4Q2...then find profit maximizing level of price and quantity


1
Expert's answer
2021-05-03T19:13:50-0400

Given:

P=10 birr

TC= 120 + 4Q2

The calculation for profit maximizing level of price and quantity is:

"=(P\times Q)-TC"

"=(10\times Q)-(120+4Q^2)"

"=10Q-120-4Q^2"

Now find derivative to maximize the profit function

"\frac {d}{dQ}(profit)=\frac{d}{dQ}(10Q-120-4Q^2)"

"10-8Q"

Now put derivative equal to zero to find the value of Q

"10-8Q=0"

"8Q=10"

"Q=1.25"

the profit maximizing quantity is 2 and the price is 10 birr